# Rethinking Pan-sharpening: Principled Design, Unified Training, and a Universal Loss Surpass Brute-Force Scaling

### v1

@@ Abstract
The field of pan-sharpening has recently seen a trend towards increasingly large and complex models, often trained on single, specific satellite datasets. This approach, however, leads to high computational overhead and poor generalization on full resolution data, a paradigm we challenge in this paper. In response to this issue, we propose PanTiny, a lightweight, single-step pan-sharpening framework designed for both efficiency and robust performance. More critically, we introduce multiple-in-one training paradigm, where a single, compact model is trained simultaneously on three distinct satellite datasets (WV2, WV3, and GF2) with different resolution and spectral information. Our experiments show that this unified training strategy not only simplifies deployment but also significantly boosts generalization on full-resolution data. Further, we introduce a universally powerful composite loss function that elevates the performance of almost all of models for pan-sharpening, pushing state-of-the-art metrics into a new era. Our PanTiny model, benefiting from these innovations, achieves a superior performance-to-efficiency balance, outperforming most larger, specialized models. Through extensive ablation studies, we validate that principled engineering in model design, training paradigms, and loss functions can surpass brute-force scaling. Our work advocates for a community-wide shift towards creating efficient, generalizable, and data-conscious models for pan-sharpening. The code is available at https://github.com/Zirconium233/PanTiny.
@@ 1 Introduction
Pan-sharpening, a fundamental image fusion task in remote sensing, aims to merge a high-resolution panchromatic (PAN) image with a lower-resolution multispectral (LRMS) image to generate a single high-resolution multispectral (HRMS) image. This fused image is crucial for numerous downstream applications, including land-cover classification, environmental monitoring, and urban planning (Masi et al. 2016; Yang et al. 2017). Early approaches were dominated by traditional methods such as Component Substitution (CS) (Carper et al. 1990; Chavez and Kwarteng 1989) and Multi-Resolution Analysis (MRA) (King and Wang 2001; Liu 2000), which, while efficient, often introduced spectral and spatial distortions due to their handcrafted prior. The advent of deep learning, particularly with models like PNN (Masi et al. 2016) and PanNet (Yang et al. 2017), revolutionized the field by learning complex mappings directly from data and significantly improving fusion quality. However, the recent pursuit of higher performance has led to a problematic trend. State-of-the-art (SOTA) methods increasingly rely on massive, complex architectures. For instance, CFDCNet (Li et al. 2025) achieves high metrics but at the cost of an astounding 55G FLOPs under 128x128 resolution images. Other models like Pan-Mamba (He et al. 2025) perform well but show limited generalization capabilities. Methods such as PanFlow (Yang et al. 2023) and diffusion-based models (Zhong et al. 2025) are often multi-step, complicating the inference process. Conversely, lightweight solutions like Pan-LUT (Cai et al. 2025), while fast, exhibit a noticeable performance gap. These specific shortcomings point to deeper, community-wide issues rooted in the pursuit of benchmark leadership over practical utility. A prevailing belief in "brute-force scaling" has led to models that are excessively large for the domain’s relatively small datasets (Deng et al. 2022). This issue is worsened by the community standard of training separate models for each satellite—a "one-dataset, one-model" philosophy that hinders both deployment efficiency and generalization. This reveals a critical issue: the majority of training for single-dataset models may contribute little to true, transferable generalization. This is evidenced by a finding in our appendix: a model fully converged on a source dataset shows nearly the same cross-domain performance as a simple baseline trained on that same source dataset for just one epoch. Secondly, the community standard of training separate models for each dataset (‘one-dataset, one-model’) hinders convenience and generalization. This is critical because true generalization is not merely about testing a single-dataset model on other datasets; as our appendix shows, such models are often just overfitting, with their cross-domain performance being matched by a model trained for just one epoch. In this work, we propose a comprehensive solution that challenges these norms. We first introduce PanTiny, a lightweight model that, AS illustrated in Figure 1 strikes a good balance between performance and efficiency, . Our extensive architectural ablations show that while scaling our model can further boost performance, it comes with diminishing returns, justifying our focus on efficiency. Second, we introduce a new “all-in-one” training paradigm in pan-sharpening domain, Through extensive experiments, we demonstrate that this approach not only simplifies the deployment pipeline but also significantly improves generalization on full-resolution data. We believe our findings support a shift towards unified models. This direction is inspired by the foundation model trend in other vision fields, suggesting a promising path for creating more generalizable pan-sharpening solutions. Finally, we design a powerful and universal composite loss function that significantly elevates the performance of all tested models, pushing the SOTA for metrics like GF2 PSNR into the 48-49 dB era. Our contributions are threefold:
@@ 2.3 Recent Advances and SOTA Models
The current landscape of pan-sharpening is dominated by advanced deep learning architectures. Inspired by successes in other vision tasks, researchers have incorporated Transformers (Zhou, Liu, and Wang 2022), State-Space Models (SSMs) like Mamba (He et al. 2025), and flow-based models (Yang et al. 2023). For instance, Pan-Mamba (He et al. 2025) leverages the efficiency of SSMs to achieve impressive results. The very recent CFDCNet (Li et al. 2025) has pushed performance metrics to new heights, but at the cost of an enormous computational load (55G FLOPs). Other approaches, such as PSCINN (Wang et al. 2024), utilize invertible neural networks to model the fusion process. While powerful, these models often come with a substantial increase in parameters and complexity. Furthermore, a common thread among these SOTA methods is their training protocol: they are almost exclusively trained and tested on a single dataset. Some works explore generalization by training on one dataset and testing on others (Chen et al. 2022), but the performance drop is often significant. The concept of an “all-in-one” model, trained jointly on multiple datasets for a single task, remains largely unexplored in pan-sharpening, representing a key opportunity that our work addresses.
@@ 3.1 Overall Architecture
Our investigation into fusion mechanisms revealed a surprising insight: in the multi-dataset training context, simplicity triumphs over complexity. We found that complex fusion strategies like cross-attention or the multi-layer “deepfusion” block from (He et al. 2025) actually degraded performance compared to a simple baseline. We attribute this to overfitting. Complex fusion modules tend to memorize dataset-specific artifacts. This "specialized knowledge" fails to generalize when the model is required to perform across multiple datasets, whereas a simpler module is forced to learn more robust, common features. The core of our network is a series of standard Transformer blocks, which effectively model long-range dependencies and perform deep feature interaction. After the Transformer body, we use a simple fusion block composed of two consecutive 3x3 convolutional layers (‘Enhanced Conv’). This design choice, validated in Table 7, proved to be the most effective and robust across all datasets.
@@ 3.2 Transformer Block
While the overall structure is inspired by the original Transformer (Vaswani et al. 2017), our implementation uses a Pre-LayerNorm (Pre-LN) configuration for improved training stability. For an input feature map $X_{l-1}$, the output $X_{l}$ of a single Transformer block is computed as:
@@
Table 1: Main quantitative comparison. All models are trained simultaneously on WV2, WV3, and GF2 datasets and evaluated on each using a single model. Our PanTiny (Big) achieves the best performance across all datasets. ‘-’ indicates a traditional, non-learning based approach. ‘nan’ indicates the model failed to train. Best results are in bold, second-best are underlined.
@@
Figure 3: Quality comparison across SoTA methods. on WV3 dataset. Refer to appendix for more results.
@@ 3.3 Loss Function
  • Focal Loss for Regression: Inspired by focal loss in classification, we adapt it for regression to prioritize "hard" pixels that are more difficult to reconstruct. Let $d_{i}=|O_{i}-G_{i}|$ be the absolute error for a given pixel $i$. Our regression-style focal loss is formulated as:
@@ 4.1 Setup
Datasets. We conduct experiments on three public datasets: WorldView-2 (WV2), WorldView-3 (WV3), and GaoFen-2 (GF2). For our primary “all-in-one” experiments, we combine the training sets of all three. We follow standard protocols for evaluation, using both reduced-resolution and full-resolution test sets. Evaluation Metrics. We provide a comprehensive evaluation using both reference and no-reference metrics. For reduced-resolution evaluation, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM) (Wang et al. 2004), Spectral Angle Mapper (SAM) (Yuhas, Goetz, and Boardman 1992), and ERGAS (Wald 2002). For full-resolution evaluation, we use the no-reference metrics Dλ, Ds, and QNR (Alparone et al. 2008). Implementation Details. Our framework is implemented in PyTorch (Paszke et al. 2019). All models are trained on a single NVIDIA RTX 4090 GPU. We use the ADAM optimizer with a learning rate of $5\times 10^{-4}$ and betas of $(0.9,0.999)$. A cosine annealing scheduler adjusts the learning rate over 500 epochs with a batch size of 16.
@@ 4.2 Quantitative Comparison and Ablations
Our experimental evaluation is designed to validate two core theses: 1) our ‘all-in-one’ training paradigm is a more robust and effective method for developing generalizable pan-sharpening models, and 2) our ‘PanTiny’ architecture achieves a superior balance of performance and efficiency compared to existing methods.
@@
Table 1 presents the main results of our study. All listed methods were trained under our unified “all-in-one” paradigm on the combined WV2, WV3, and GF2 datasets, and evaluated on each one’s test set using a single set of model weights. We present two versions of our model: PanTiny (Small), our ultra-lightweight variant, and PanTiny (Big), our primary model that achieves the best performance. We include both to highlight the excellent efficiency of our base architecture and the SOTA performance achieved with a modest increase in size. Our proposed PanTiny (Big) achieves SOTA performance across all three datasets, outperforming both classic and recent methods. Notably, it surpasses CFDCNet (Li et al. 2025), a much larger model, on all metrics. It also significantly outperforms other lightweight methods like PanFlow (Yang et al. 2023). The results for PSCINN (Wang et al. 2024) highlight their instability in a multi-dataset setting, as it failed to complete training, further validating our design choices for robustness.
@@
A core contribution of our work is the “all-in-one” training paradigm. Prior work on generalization often involves training on a single dataset and testing on others. However, due to the significant domain gap between satellite datasets, this approach struggles to produce a truly universal model. Our preliminary tests (detailed in the appendix) show that a model trained for just one epoch on a source dataset can sometimes match the cross-dataset performance of a fully-trained so-called "general" model, suggesting the latter may be overfitting. Our “all-in-one” approach directly addresses this by exposing the model to multiple domains during training. As shown in Table 3, this has a profound effect. Complex models like Pan-Mamba and PSCINN see a significant performance drop compared to their specialized, separately trained counterparts. In contrast, our ‘PanTiny (Big)’ model is remarkably robust, with only a minor drop of  0.3 PSNR on WV2/GF2 and almost no change on WV3. Furthermore, Table 4 demonstrates that this paradigm enhances generalization on real-world, full-resolution data. For all tested models, switching from separate to all-in-one training results in a substantial improvement in the no-reference QNR metric on the WV2 full-resolution dataset. While PanFlow achieves the highest QNR, the universal improvement underscores the power of our proposed training method as a general tool for improving pan-sharpening model generalization. Note that some models like Pan-Mamba and CFDCNet are excluded as their codebase does not support variable inference resolutions without modification; details are in the supplement.
@@
To demonstrate the superiority of our network architecture, we conducted a benchmark where all competing methods were trained with our proposed composite loss. As shown in Table 5, even when other methods benefit from our improved training process, our ‘PanTiny (Big)’ model still secures a top-tier position, surpassed only by the much larger CFDCNet (+1.6M params, +53G FLOPs). Our exploration process, also detailed in the table, further justifies our choices:
  • Limitations of Naive Scaling: We first explored two intuitive designs: ‘DeepPNN’, a deeper and wider version of PNN, and ‘ResAtten’, which combines a standard ResNet backbone with attention. While both achieve competitive performance (with ResAtten reaching the highest PSNR on WV2), they require a significantly larger number of parameters (over 260K). This demonstrates that simply scaling up or using generic vision backbones is not the most efficient path to SOTA performance.
@@
The effect of scaling is further explored in Table 6. By expanding our model to a ‘Huge’ version (196K params), we can nearly match the performance of CFDCNet on GF2 (48.85 vs 49.07 PSNR), but at less than 12% of its parameter count. This reinforces our core argument against inefficient scaling and validates our choice of ‘PanTiny (Big)’as the optimal model.
@@
We ablate the fusion and refinement blocks in Table 7 and Table 8. For fusion, our simple “Enhanced Conv” (2-layer Conv) outperforms both a simpler “1x1 Conv” and more complex attention-based mechanisms. Notably, “DeepFusion” from Pan-Mamba, which is highly effective in a single-dataset setting, performs poorly here, again suggesting it overfits. For refinement, a simple “Conv” layer is optimal. Adding complexity via a larger convolution or channel attention increases parameters without a consistent performance benefit, and in some cases, hurts the results.
Table 7: Ablation on the fusion module. (Showing only PSNR).
@@
Table 8: Ablation on the refinement module. (Showing only PSNR).
@@
The significant performance boost of our method stems not only from its architecture but also from our carefully engineered loss function—a key innovation in its own right. As shown in Table 5, our composite loss provides a universal performance uplift to all tested models, demonstrating its power as a general tool for the community. Our extensive search for the optimal configuration, detailed in the appendix, began with evaluating individual components. This revealed that SSIM loss is particularly effective, especially for the GF2 dataset. Building on this insight, we found that assigning a high weight to the SSIM component consistently yielded improvements. Our final weights ($\lambda_{1}=1.5,\lambda_{2}=4.0,\lambda_{3}=1.5$) represent the best-performing combination from this exhaustive search. This well-tuned, composite loss has proven to be a cornerstone of our work, enabling a significant leap in performance and pushing the pan-sharpening field into a new era of 48-49 PSNR on the GF2 dataset.

### v2

@@ Abstract
The field of pan-sharpening has recently seen a trend towards increasingly large and complex models, often trained on single, specific satellite datasets. This approach, however, leads to high computational overhead and poor generalization on full resolution data, a paradigm we challenge in this paper. In response to this issue, we propose PanTiny, a lightweight, single-step pan-sharpening framework designed for both efficiency and robust performance. More critically, we introduce multiple-in-one training paradigm, where a single, compact model is trained simultaneously on three distinct satellite datasets (WV2, WV3, and GF2) with different resolution and spectral information. Our experiments show that this unified training strategy not only simplifies deployment but also significantly boosts generalization on full-resolution data. Further, we introduce a universally powerful composite loss function that elevates the performance of almost all of models for pan-sharpening, pushing state-of-the-art metrics into a new era. Our PanTiny model, benefiting from these innovations, achieves a superior performance-to-efficiency balance, outperforming most larger, specialized models. Through extensive ablation studies, we validate that principled engineering in model design, training paradigms, and loss functions can surpass brute-force scaling. Our work advocates for a community-wide shift towards creating efficient, generalizable, and data-conscious models for pan-sharpening. The code is available at https://github.com/Zirconium233/PanTiny
@@ 1 Introduction
Pan-sharpening, a fundamental image fusion task in remote sensing, aims to merge a high-resolution panchromatic (PAN) image with a lower-resolution multispectral (LRMS) image to generate a single high-resolution multispectral (HRMS) image. This fused image is crucial for numerous downstream applications, including land-cover classification, environmental monitoring, and urban planning (Masi et al. 2016; Yang et al. 2017). Early approaches were dominated by traditional methods such as Component Substitution (CS) (Carper et al. 1990; Chavez and Kwarteng 1989) and Multi-Resolution Analysis (MRA) (King and Wang 2001; Liu 2000), which, while efficient, often introduced spectral and spatial distortions due to their handcrafted prior. The advent of deep learning, particularly with models like PNN (Masi et al. 2016) and PanNet (Yang et al. 2017), revolutionized the field by learning complex mappings directly from data and significantly improving fusion quality.
However, the recent pursuit of higher performance has led to a problematic trend. State-of-the-art (SOTA) methods increasingly rely on massive, complex architectures. For instance, CFDCNet (Li et al. 2025) achieves high metrics but at the cost of an astounding 55G FLOPs under $\times 128$ 128 resolution images. Other models like Pan-Mamba (He et al. 2025) perform well but show limited generalization capabilities. Methods such as PanFlow (Yang et al. 2023) and diffusion-based models (Zhong et al. 2025) are often multi-step, complicating the inference process. Conversely, lightweight solutions like Pan-LUT (Cai et al. 2025), while fast, exhibit a noticeable performance gap.
These specific shortcomings point to deeper, community-wide issues rooted in the pursuit of benchmark leadership over practical utility. A prevailing belief in “brute-force scaling” has led to models that are excessively large for the domain’s relatively small datasets (Deng et al. 2022). This issue is worsened by the community standard of training separate models for each satellite—a “one-dataset, one-model” philosophy that hinders both deployment efficiency and generalization. This reveals a critical issue: the majority of training for single-dataset models may contribute little to true, transferable generalization. This is evidenced by a finding in our appendix: a model fully converged on a source dataset shows nearly the same cross-domain performance as a simple baseline trained on that same source dataset for just one epoch.
In this work, we propose a comprehensive solution that challenges these norms. We first introduce PanTiny, a lightweight model that, as illustrated in Figure 1, strikes a good balance between performance and efficiency. Our extensive architectural ablations show that while scaling our model can further boost performance, it comes with diminishing returns, justifying our focus on efficiency. Second, we introduce a new “all-in-one” training paradigm in pan-sharpening domain, through extensive experiments, we demonstrate that this approach not only simplifies the deployment pipeline but also significantly improves generalization on full-resolution data. We believe our findings support a shift towards unified models. This direction is inspired by the foundation model trend in other vision fields, suggesting a promising path for creating more generalizable pan-sharpening solutions. Finally, we design a powerful and universal composite loss function that significantly elevates the performance of all tested models, pushing the SOTA for metrics like GF2 PSNR into the 48-49 dB era.
Our contributions are threefold:
@@ 2.3 Recent Advances and SOTA Models
The current landscape of pan-sharpening is dominated by advanced deep learning architectures. Inspired by successes in other vision tasks, researchers have incorporated Transformers, State-Space Models (SSMs) like Mamba (He et al. 2025), and flow-based models (Yang et al. 2023). For instance, Pan-Mamba leverages the efficiency of SSMs to achieve impressive results. The very recent CFDCNet (Li et al. 2025) has pushed performance metrics to new heights, but at the cost of an enormous computational load (55G FLOPs). Other approaches, such as PSCINN (Wang et al. 2024), utilize invertible neural networks to model the fusion process. While powerful, these models often come with a substantial increase in parameters and complexity. Furthermore, a common thread among these SOTA methods is their training protocol: they are almost exclusively trained and tested on a single dataset. Some works explore generalization by training on one dataset and testing on others (Zhang et al. 2024), but the performance drop is often significant. The concept of an “all-in-one” model, trained jointly on multiple datasets for a single task, remains largely unexplored in pan-sharpening, representing a key opportunity that our work addresses.
@@ 3.1 Overall Architecture
Our investigation into fusion mechanisms revealed a surprising insight: in the multi-dataset training context, simplicity triumphs over complexity. We found that complex fusion strategies like cross-attention or the multi-layer “deepfusion” block from actually degraded performance compared to a simple baseline. We attribute this to overfitting. Complex fusion modules tend to memorize dataset-specific artifacts. This “specialized knowledge” fails to generalize when the model is required to perform across multiple datasets, whereas a simpler module is forced to learn more robust, common features. The core of our network is a series of standard Transformer blocks, which effectively model long-range dependencies and perform deep feature interaction. After the Transformer body, we use a simple fusion block composed of two consecutive 3x3 convolutional layers (“Enhanced Conv”). This design choice, validated in Table 7, proved to be the most effective and robust across all datasets.
@@ 3.2 Transformer Block
While the overall structure is inspired by the original Transformer, our implementation uses a Pre-LayerNorm (Pre-LN) configuration for improved training stability. For an input feature map $X_{l-1}$, the output $X_{l}$ of a single Transformer block is computed as:
@@
Table 1: Main quantitative comparison. All models are trained simultaneously on WV2, WV3, and GF2 datasets and evaluated on each using a single model. Our PanTiny (Big) achieves the best performance across all datasets. ‘-’ indicates a traditional, non-learning based approach. Best results are in bold, second-best are underlined.
Figure 3: Quality comparison across SOTA methods on WV3 dataset. Refer to appendixs for more results.
@@ 3.3 Loss Function
  • Focal Loss for Regression: Inspired by focal loss in classification, we adapt it for regression to prioritize “hard” pixels that are more difficult to reconstruct. Let $d_{i}=|O_{i}-G_{i}|$ be the absolute error for a given pixel $i$. Our regression-style focal loss is formulated as:
@@ 4.1 Setup
Datasets. We conduct experiments on three public datasets: WorldView-2 (WV2), WorldView-3 (WV3), and GaoFen-2 (GF2). For our primary “all-in-one” experiments, we combine the training sets of all three. We follow standard protocols for evaluation, using both reduced-resolution and full-resolution test sets.
Evaluation Metrics. We provide a comprehensive evaluation using both reference and no-reference metrics. For reduced-resolution evaluation, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM) (Wang et al. 2004), Spectral Angle Mapper (SAM) (Yuhas, Goetz, and Boardman 1992), and ERGAS (Wald 2002). For full-resolution evaluation, we use the no-reference metrics Dλ, Ds, and QNR (Alparone et al. 2008).
Implementation Details. Our framework is implemented in PyTorch. All models are trained on a single NVIDIA RTX 4090 GPU. We use the ADAM optimizer with a learning rate of $5\times 10^{-4}$ and betas of (0.9,0.999)(0.9,0.999)( 0.9 , 0.999 ). A cosine annealing scheduler adjusts the learning rate over 500 epochs with a batch size of 16.
@@ 4.2 Quantitative Comparison and Ablations
Our experimental evaluation is designed to validate two core theses: 1) our “all-in-one” training paradigm is a more robust and effective method for developing generalizable pan-sharpening models, and 2) our “PanTiny” architecture achieves a superior balance of performance and efficiency compared to existing methods.
@@
Table 1 presents the main results of our study. All listed methods were trained under our unified “all-in-one” paradigm on the combined WV2, WV3, and GF2 datasets, and evaluated on each one’s test set using a single set of model weights. We present two versions of our model: PanTiny (Small), our ultra-lightweight variant, and PanTiny (Big), our primary model that achieves the best performance. We include both to highlight the excellent efficiency of our base architecture and the SOTA performance achieved with a modest increase in size.
Our proposed PanTiny (Big) achieves SOTA performance across all three datasets, outperforming both classic and recent methods. Notably, it surpasses CFDCNet (Li et al. 2025), a much larger model, on all metrics. It also significantly outperforms other lightweight methods like PanFlow (Yang et al. 2023). The results for PSCINN (Wang et al. 2024) highlight their instability in a multi-dataset setting, as it failed to complete training, further validating our design choices for robustness.
@@
A core contribution of our work is the “all-in-one” training paradigm. Prior work on generalization often involves training on a single dataset and testing on others. However, due to the significant domain gap between satellite datasets, this approach struggles to produce a truly universal model. Our preliminary tests (detailed in the appendix) show that a model trained for just one epoch on a source dataset can sometimes match the cross-dataset performance of a fully-trained so-called “general” model, suggesting the latter may be overfitting.
Our “all-in-one” approach directly addresses this by exposing the model to multiple domains during training. As shown in Table 3, this has a profound effect. Complex models like Pan-Mamba and PSCINN see a significant performance drop compared to their specialized, separately trained counterparts. In contrast, our “PanTiny (Big)” model is remarkably robust, with only a minor drop of  0.3 PSNR on WV2/GF2 and almost no change on WV3.
Furthermore, Table 4 demonstrates that this paradigm enhances generalization on real-world, full-resolution data. For all tested models, switching from separate to all-in-one training results in a substantial improvement in the no-reference QNR metric on the WV2 full-resolution dataset. While PanFlow achieves the highest QNR, the universal improvement underscores the power of our proposed training method as a general tool for improving pan-sharpening model generalization. Note that some models like Pan-Mamba and CFDCNet are excluded as their codebase does not support variable inference resolutions without modification; details are in the supplement.
@@
To demonstrate the superiority of our network architecture, we conducted a benchmark where all competing methods were trained with our proposed composite loss. As shown in Table 5, even when other methods benefit from our improved training process, our “PanTiny (Big)” model still secures a top-tier position, surpassed only by the much larger CFDCNet (+1.6M params, +53G FLOPs). Our exploration process, also detailed in the table, further justifies our choices:
  • Limitations of Naive Scaling: We first explored two intuitive designs: “DeepPNN”, a deeper and wider version of PNN, and “ResAtten”, which combines a standard ResNet backbone with attention. While both achieve competitive performance (with ResAtten reaching the highest PSNR on WV2), they require a significantly larger number of parameters (over 260K). This demonstrates that simply scaling up or using generic vision backbones is not the most efficient path to SOTA performance.
@@
The effect of scaling is further explored in Table 6. By expanding our model to a ‘Huge’ version (196K params), we can nearly match the performance of CFDCNet on GF2 (48.85 vs 49.07 PSNR), but at less than 12% of its parameter count. This reinforces our core argument against inefficient scaling and validates our choice of “PanTiny (Big)” as the optimal model.
@@
We ablate the fusion and refinement blocks in Table 7 and Table 8. For fusion, our simple “Enhanced Conv” (2-layer Conv) outperforms both a simpler “$\times 1$ 1 Conv” and more complex attention-based mechanisms. Notably, “DeepFusion” from Pan-Mamba, which is highly effective in a single-dataset setting, performs poorly here, again suggesting it overfits. For refinement, a simple “Conv” layer is optimal. Adding complexity via a larger convolution or channel attention increases parameters without a consistent performance benefit, and in some cases, hurts the results.
@@
Table 7: Ablation on the fusion module. (Showing only PSNR).
@@
Table 8: Ablation on the refinement module. (Showing only PSNR).
@@
The significant performance boost of our method stems not only from its architecture but also from our carefully engineered loss function—a key innovation in its own right. As shown in Table 5, our composite loss provides a universal performance uplift to all tested models, demonstrating its power as a general tool for the community. Our extensive search for the optimal configuration, detailed in the appendixs, began with evaluating individual components. This revealed that SSIM loss is particularly effective, especially for the GF2 dataset. Building on this insight, we found that assigning a high weight to the SSIM component consistently yielded improvements. Our final weights ($\lambda_{1}=1.5,\lambda_{2}=4.0,\lambda_{3}=1.5$) represent the best-performing combination from this exhaustive search. This well-tuned, composite loss has proven to be a cornerstone of our work, enabling a significant leap in performance and pushing the pan-sharpening field into a new era of 48-49 PSNR on the GF2 dataset.